Two syringes of different cross sections (without needles) filled with water are connected with a tightly fitted rubber tube filled with water. Diameters of the smaller piston and larger piston are 1.0cm and 3.0 cm respectively. If the smaller piston is pushed in through 6.0 cm, how much does the larger piston move out
1
Expert's answer
2018-12-20T09:19:10-0500

Since the water volume through pistons motion must be constant, we obtain

(πd_1^2)/4 h_1=(πd_2^2)/4 h_2

So

h_2=(h_1 d_1^2)/(d_2^2 )=〖6.0×1.0〗^2/〖3.0〗^2 =2/3 cm=0.67 cm